OpenStudy (anonymous):

what multiplies to give you -24 and adds to give you -9?

zepdrix (zepdrix):

So let's start with factors of 24. 1*24 2*12 3*8 4*6 Hmmm I don't think any of those will give you -9. -8+3=-5 -12+2=-10 I think those are the closest ones... maybe I missed one somewhere XD There don't appear to be any integers that will get you there.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

there is no #, must be a trick question

OpenStudy (anonymous):

xy=-24 x+y=-9 so: x=-9-y substituting: -9y-y^2=-24 or y^2+9y-24=0 x=[-9+-sqrt(81+96)]/2. So it is not a natural number

OpenStudy (anonymous):

y=-48/[-9+-sqrt(81+96)]